OpenStudy (anonymous):

find the difference quotient, [f(X+h)-f(x)]/h, for the function and simplify it. P(X)=-10x^3+2

OpenStudy (freckles):

I guess P=f

OpenStudy (freckles):

do you know how to find P(x+h)?

OpenStudy (freckles):

P(x)=-10x^3+2 so to find P(x+h) you replace the x's in P(x)=-10x^3+2 with (x+h)'

OpenStudy (anonymous):

ok

OpenStudy (freckles):

so what is P(x+h)=?

OpenStudy (anonymous):

p(x+h)

OpenStudy (freckles):

what is that equal to given p(x)=-10x^3+h

OpenStudy (freckles):

oops p(x)=-10x^3+2*

OpenStudy (freckles):

\[\text{ if } P(x)=-10x^3+2 \text{ then } \\ P(fish)=-10(fish)^3+2 \\ \text{ so } P(x+h)=?\]

OpenStudy (anonymous):

should i expand (X+h)^3

OpenStudy (freckles):

eventually need to simplify*

OpenStudy (freckles):

but do you think you know what P(x+h) is and if so what is it

OpenStudy (anonymous):

im at [(-10x-10)^3+10x^3-2]/h

OpenStudy (anonymous):

thats -10(x+h)^3

OpenStudy (freckles):

well P(x+h) isn't -10(x+h)^3 but it is -10(x+h)^3+2 remember P(x)=-10x^3+2

OpenStudy (freckles):

\[\frac{p(x+h)-p(x)}{h} \\ =\frac{[-10(x+h)^3+2]-[-10x^3+2]}{h}\]

OpenStudy (anonymous):

haha yup still doesnt dodisturb:)

OpenStudy (freckles):

no distribute the minus sign in front of the second grouping there

OpenStudy (freckles):

you will see something cancel after distributing the minus in front of the second grouping

OpenStudy (freckles):

\[\frac{p(x+h)-p(x)}{h} \\ =\frac{[-10(x+h)^3+2] \color{red}{-}[-10x^3+2]}{h}\] see that red - in front of the [ ]

OpenStudy (freckles):

-(a+b)=-a-b

OpenStudy (freckles):

it changes the sign of the terms inside the [ ]

OpenStudy (anonymous):

yes i know lolololol

OpenStudy (freckles):

do you see something cancel then?

OpenStudy (anonymous):

how would i simplify it

OpenStudy (anonymous):

yea the 2 lol i told you it didnt matter lol

OpenStudy (freckles):

well that 2 does matter from the p(x+h) so it cancel with the 2 from the p(x)

OpenStudy (freckles):

if you don't see that p(x+h)=-10(x+h)^3+2 then both the 2's would have not cancel

OpenStudy (anonymous):

can just like finish the problem ill analyze it your too good of a teacher:) or at least tell me what to do with the (x+h)^3

OpenStudy (anonymous):

do i have to expand it????

OpenStudy (freckles):

\[\frac{p(x+h)-p(x)}{h} \\ =\frac{[-10(x+h)^3+2]-[-10x^3+2]}{h} = \frac{-10(x+h)^3+2+10x^3-2}{h}\] now you cancel the 2's and as suggested earlier by you need to exapnd the (x+h)^3

OpenStudy (freckles):

after you expand the (x+h)^3 you will see another pair cancel

OpenStudy (freckles):

what do you have after doing that?

OpenStudy (anonymous):

[-10x^2-10h^3-30x^2h-30xh^2+10x^2]/h

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

sorry this is the part ive been stuck on since the beggining

OpenStudy (freckles):

\[(x+h)^3=x^3+3x^2h+3xh^2+h^3 \\ -10(x+h)^3=-10x^3-30x^2h-30xh^2-10h^3 \] when did you get -10x^2 +10x^2 ?

OpenStudy (freckles):

\[\frac{p(x+h)-p(x)}{h} \\ =\frac{[-10(x+h)^3+2]-[-10x^3+2]}{h} = \frac{-10(x+h)^3+2+10x^3-2}{h} \\ \frac{-10x^3-30x^2h-30xh^2-10h^3+10x^3}{h}\] you should see that -10x^3+10x^3=?

OpenStudy (anonymous):

cancell

OpenStudy (freckles):

yes not the remaining terms on top have an h and there is also a h on bottom

OpenStudy (freckles):

so i will let you decide what do we need to do next

OpenStudy (freckles):

now*

OpenStudy (anonymous):

factor it out

OpenStudy (freckles):

yes factor a h on top and then cancel that h with the h on bottom since h/h=1

OpenStudy (anonymous):

thank you<333